Grade 9Physical Chemistry

Calculate the number of the constituents atom in 53g of Na2CO3

PrathMesh lande
9 Years agoGrade 9

1 Answer

Me
9 Years ago
Given mass=53g
Molar mass of Na2CO3 = 23×2+12+48 = 46+60= 106g
No. Of moles = given mass /molar mass
                        = 53/106
                        =0.5
Now 1 mole = 6.02×10^23 atoms
So, 0.5 moles = 0.5×6.02×10^23. = 3.010×10^23 atoms